OpenStudy (anonymous):

F(x)=3x^2 - x - 2 if F(x)= -2, what is X

OpenStudy (anonymous):

F(x) = -2 when x = 0. Setting x=0 gives the y-intercept for the graph of f(x).

OpenStudy (anonymous):

is there another answer besides zero?

OpenStudy (anonymous):

lol just plug in ur point as y :) -2=3x^2 - x - 2

OpenStudy (anonymous):

can u do it or do u want the answer and for Tims, y-intercept does not equal solution

OpenStudy (anonymous):

x=0 and x=1/3 :)

OpenStudy (anonymous):

well, i see zero is an answer after plugging in, but apparently there is another answer that = -2 as well. and i cant find it

OpenStudy (anonymous):

its 1/3 newkidoncampus :P

OpenStudy (anonymous):

haha. ok i tried 1/2 but not 1/3. thank you.

OpenStudy (anonymous):

-2=3x^2 - x - 2 0=3x^2-x 0=x(3x-1) x=0 and 3x-1=0 x=1/3 :)

OpenStudy (anonymous):

this is how u do it :)

OpenStudy (anonymous):

ohh ok. got it. ;)